OpenStudy (anonymous):

Solve the system algebraically 2x+y-10=0 X-y-4=0 What is the value of y? A) 1/3 B) 2/3 C) 14/3

OpenStudy (anonymous):

2x+y-10 = 0 x - y - 4 = 0 =========== 3x - 14 = 0 3x = 14 x = 14/3 x - y = 4 14/3 - y = 4 14/3 - 4 = y 14/3 - 12/3 = y 2/3 = y y = 2/3 or Multiply the second equation by 2 and then subtract the new one from the first one. We get 3y -2 = 0, so y = 2/3. Hence the value of y is 2/3. Depends which way you like to solve it...